# Rob A. J. Eyles-Ferris,1 Andrew King,1,2,3 Rhaana L. C. Starling,1 Peter G. Jonker,4 Andrew J. Levan,4 Antonio Martin-Carrillo,5 Tanmoy Laskar,6,4 Jillian C. Rastinejad,7,8 Nikhil Sarin,9,10 Nial R. Tanvir,1 Benjamin P. Gompertz,11,12 Nusrin Habeeb,1 Paul T. O’Brien1 and Massimiliano De Pasquale13

### v1

@@ 3.4 Formation of the required system
While the system we have described can produce the electromagnetic signal observed in GRB 250702B, we should briefly consider the likelihood of it arising. As mentioned in Section 3.1, the extremely bound nature of the WD’s orbit may be rare. From Hills (1988) and Cufari et al. (2022), the analytically predicted eccentricity of a bound star following the Hills mechanism is given by
$e\simeq 1-\frac{2}{\beta}\left(\frac{M_{\bullet}}{M_{\star}}\right)^{-1/3},$ (2)
where $\beta=r_{t,{\rm binary}}/r_{p}$ and $M_{\star}$ is the mass of the primary which we assume to be our WD i.e. $0.6~M_{\odot}$ [^8] This is also a reasonable assumption given the mass ratios of WD binaries observed in the Milky Way (e.g. Munday et al., 2024).. Equation 2 gives $e\sim 0.92$ for our assumed system assuming $\beta=1$ and Cufari et al. (2022) show that numerical methods indicate the peak of the probability density function of $e$ to be roughly consistent, although slightly higher, than the predictions of Equation 2. Our inferred orbit therefore seems reasonable assuming a suitable encounter between a WD binary and an IMBH occurs.
@@
This latter point is difficult to quantify due to the poorly constrained rates of both IMBHs and WD binaries. If we assume a limiting flux for Fermi-GBM of $5\times 10^{-8}$ erg cm-2 s-1 (Liu et al., 2025, see also Perley et al. (2016)), then triggers as luminous as GRB 250702B’s peak should be detectable out to a comoving distance of $\sim 24.6$ Gpc. With only this one source uniquely observed over the 17 years of Fermi’s mission, this suggests a volumetric rate of $\sim 10^{-6}$ Gpc-3 yr-1, consistent with the estimates of Gompertz et al. (in prep.). From the other perspective, the assumed rate of typical TDEs ($10^{-5}-10^{-4}$ galaxy-1 yr-1, see Gezari, 2021, and references therein) represents an upper limit on events like this given the relative rates of observed IMBH TDEs [^9] At time of writing, $\sim 200$ to 300 TDE candidates are known of which two to three have been linked to IMBHs.. The disruption of a WD is likely even rarer and there are no confirmed candidates observationally. However, Ye et al. (2023) predict the capture rate of WDs by nuclear IMBHs to be of order $\sim 10^{-6}$ galaxy-1 yr-1, consistent with estimates from MacLeod et al. (2014, see also () and references therein). For IMBHs in the outskirts of a galaxy, this is likely to be several orders of magnitude smaller especially given the eccentricity and beaming requirements we derive above. This implies a volumetric rate of order $10^{-3}$ Gpc-3 yr-1 and broadly consistent with the rarity implied by our calculated volumetric rate.
4 Conclusion
@@
To conclude, we have presented and analysed the results of X-ray observations of GRB 250702B and have also identified several key features from across the electromagnetic spectrum that must be explained in any model of this event. We have shown that all of these features can, withe data currently available to us, be exhibited by the tidal disruption of white dwarf by an intermediate mass black hole. These features and how they can arise from a WD-IMBH TDE are summarised below:
@@ Acknowledgements
The data used in this work are publicly available, either from the UKSSDC or the Chandra Data Archive (see footnotes above for links).

### v2

@@ 3.4 Formation of the required system
While the system we have described can produce the electromagnetic signal observed in GRB 250702B, we should briefly consider the likelihood of it arising. As mentioned in Section 3.1, the extremely bound nature of the WD’s orbit may be rare. From Hills88 and Cufari22, the analytically predicted eccentricity of a bound star following the Hills mechanism is given by
$e\simeq 1-\frac{2}{\beta}\left(\frac{M_{\bullet}}{M_{\rm WD}}\right)^{-1/3},$ (4)
where $\beta=r_{t,{\rm binary}}/r_{p}$ and $M_{\rm WD}$ is the mass of the primary which we assume to be our WD. For our MCMC chain, assuming $\beta=1$, Equation 4 predicts eccentricities marginally higher but broadly consistent with the fitted values. Cufari22 show that numerical methods indicate the peak of the probability density function of $e$ to be roughly consistent, although slightly higher, than the predictions of Equation 4 and therefore our inferred orbits are reasonable assuming a suitable encounter between a WD binary and an IMBH occurs. These orbital properties will result in a fast runaway from initial binding to a final disruption implying the first stripping event occurred only shortly before disruption. As gravitational wave emission will change the orbit only very slowly, the initial orbit must also be similar to this final orbit. Together, these imply the WD was bound to the IMBH within a few tens of orbits of the final disruption.
@@
It is difficult to quantify how likely such the encounter between a WD binary and an IMBH will be due to the poorly constrained rates of both. If we assume a limiting flux for Fermi-GBM of $5\times 10^{-8}$ erg cm-2 s-1 (Liu25, see also Perley16), then triggers as luminous as GRB 250702B’s peak should be detectable out to a luminosity distance of $\sim 53.1$ Gpc. Accounting for time dilation effects, with only this one source uniquely observed over the 17 years of Fermi’s mission, this suggests a volumetric rate of $\sim 10^{-4}$ Gpc-3 yr-1, consistent with the estimates of Gompertz25. From the other perspective, the assumed rate of typical TDEs ($10^{-5}-10^{-4}$ galaxy-1 yr-1, see Gezari21, and references therein) represents an upper limit on events like this given the relative rates of observed IMBH TDEs [^11] At time of writing, $\sim 200$ to 300 TDE candidates are known of which two to three have been linked to IMBHs.. The disruption of a WD is likely even rarer and there are no confirmed candidates observationally. However, Ye23 predict the capture rate of WDs by nuclear IMBHs to be of order $\sim 10^{-6}$ galaxy-1 yr-1, consistent with estimates from Macleod14. For IMBHs in the outskirts of a galaxy, this is likely to be several orders of magnitude smaller especially given the eccentricity and beaming requirements we derive above. This implies a volumetric rate of order $10^{-3}$ Gpc-3 yr-1 and consistent with the rarity implied by our calculated volumetric rate.
4 Conclusions
@@
We have presented and analysed the results of X-ray observations of GRB 250702B and have also identified several key features from across the electromagnetic spectrum that must be explained in any model of this event. We have shown that all of these features can be exhibited by the tidal disruption of white dwarf by an intermediate mass black hole. These features and how they can arise from a WD-IMBH TDE are summarised below:
@@ Acknowledgements
We thank the referee for their insightful and thorough comments that have significantly improved this manuscript.
@@
The data used in this work are publicly available, either from the UKSSDC or the Chandra Data Archive (see footnotes above for links), or Li25.